### v1

# PRISM [Uncaptioned image] of Opinions: A Persona‑Reasoned Multimodal Framework for User-centric Conversational Stance Detection

Abstract
The rapid proliferation of multimodal social media content has driven research in Multimodal Conversational Stance Detection (MCSD), which aims to interpret users’ attitudes toward specific targets within complex discussions. However, existing studies remain limited by: 1) pseudo-multimodality, where visual cues appear only in source posts while comments are treated as text-only, misaligning with real-world multimodal interactions; and 2) user homogeneity, where diverse users are treated uniformly, neglecting personal traits that shape stance expression. To address these issues, we introduce U-MStance, the first user‑centric MCSD dataset, containing over 40k annotated comments across six real-world targets. We further propose PRISM, a Persona-Reasoned multImodal Stance Model for MCSD. PRISM first derives longitudinal user personas from historical posts and comments to capture individual traits, then aligns textual and visual cues within conversational context via Chain‑of‑Thought to bridge semantic and pragmatic gaps across modalities. Finally, a mutual task reinforcement mechanism is employed to jointly optimize stance detection and stance-aware response generation for bidirectional knowledge transfer. Experiments on U‑MStance demonstrate that PRISM yields significant gains over strong baselines, underscoring the effectiveness of user-centric and context-grounded multimodal reasoning for realistic stance understanding.
1 Introduction
Social media have become a dominant platform for the expression of opinions, making their analysis crucial to understand divisive societal topics such as politics, public health, and social justice [26-glandt2021stance, 27-liang2022zero]. Stance detection, the task of identifying a user’s position toward a target, has become vital for charting public discourse and polarization. As online interactions grow more complex, the field has evolved from simple textual analysis to modeling rich, multi-party conversational contexts [28-li2023improved, 29-li2023contextual, 30-niu2024challenge].
Figure 1: Comparison of a conventional approach for MCSD (left) with our proposed PRISM framework (right).
The recent shift towards Multimodal Conversational Stance Detection (MCSD), exemplified by the MmMtCSD dataset [1-niu2024multimodal], marks a crucial step forward. However, despite this progress, two fundamental limitations persist across the field, hindering authentic stance modeling. The first is a pseudo-multimodality, where visual cues appear only in source posts while comments remain text‑only, diverging from the multimodal nature of real-world interactions. The second is user homogeneity, where diverse individuals are modeled uniformly, ignoring personal traits that shape stance expression. These limitations can be clearly observed in the example shown in Figure 1. User Carrots combines an image and textual content to contrast Obama and Trump, conveying a derogatory attitude toward Trump, which highlights the necessity of multimodal cues for accurate stance understanding. Meanwhile, Zeke expresses sympathy for Trump, and Bgad challenges this view, revealing conflicting stances. A model that remains blind to their individual personas would fail to grasp the reason for their disagreement, leading to incorrect stance interpretations.
To catalyze progress beyond these limitations, we therefore introduce U-MStance, the first dataset designed for user-centric multimodal conversational stance detection. Comprising over 40k annotated comments across six real-world targets, U-MStance uniquely incorporates multimodality in both source posts and conversational replies, providing a realistic and challenging benchmark that captures the nuances of user-centric, multimodal interactions. Building upon this critical new dataset, we further propose PRISM, a Persona-Reasoned multImodal Stance Model for MCSD. Specifically, PRISM first explores the underlying personal tendencies behind stance expressions and derives longitudinal user persona traits inspired by Big Five (OCEAN) theory [44-jiang2024personallm] from historical posts and comments. Then, it introduces a rationalized cross-modal grounding module, leveraging Chain-of-Thought (CoT) reasoning to explicitly bridge the profound semantic and pragmatic gap between modalities in conversational contexts. Ultimately, these personas and contextual representations are fed into a mutual task reinforcement mechanism, which jointly optimizes stance classification and stance-aware response generation, enabling bidirectional knowledge transfer that enhances both detection performance and the modeling of user communication styles. Comprehensive experiments on our U-MStance dataset show that PRISM achieves significant improvements over strong baselines across multiple metrics. The main contributions can be summarized as follows:
  • To our knowledge, U-MStance is the first user-centric multimodal conversational stance detection dataset, which extends multimodality beyond posts to comments, and incorporates user information for more realistic and challenging stance understanding.
  • We propose PRISM, a framework that models user personas and deepens cross-modal correlations for improved user-centric stance detection.
  • Extensive experiments conducted on our U-MStance dataset demonstrate that PRISM significantly outperforms strong baselines, confirming both its effectiveness and generalization ability.
2 Related Works
2.1 Stance Datasets
The development of stance detection datasets has evolved from analyzing isolated texts to modeling complex interactions. Early benchmarks like SemEval-2016 Task 6 [2-mohammad2016semeval] and VAST [3-allaway2020zero] focused on sentence-level textual stance detection, establishing foundational models but overlooking the context of dialogue. This limitation spurred the creation of conversational datasets such as SRQ [4-villa2020stance] and CANT-CSD [5-li2023improved], which treated stance as a dynamic process unfolding within a conversation. The most recent advancement is the shift toward multimodal stance detection, which incorporates visual elements alongside text to better reflect modern social media discourse. Despite these advances, even the latest datasets for MCSD, such as MmMtCSD [1-niu2024multimodal], still face two key shortcomings that hinder realistic stance modeling. First, multimodality is handled only superficially, with source posts treated as multimodal but conversational replies reduced to text-only. In addition, information about users, including their posting history, is largely ignored, overlooking contextual factors that shape an individual’s expressed stance and leaving a substantial gap in building comprehensive and authentic stance detection models.
2.2 Stance Detection Methods
Methodological advancements in stance detection has spanned multiple task settings in-target [6-li2021multi], cross-target [7-ding2024cross, 8-liang2021target, 9-wei2019modeling], and zero-shot scenarios [10-allaway2020zero, 11-liang2022zero, 12-zhu2022enhancing]. Early deep learning approaches primarily fine-tuned pre-trained models such as BERT [13-cambria2018senticnet]. However, the field was revolutionized by the emergence of Large Language Models (LLMs), whose powerful semantic understanding and adaptability introduced new capabilities [14-cheng2024emotion, 15-cheng2024mips, 16-li2024mitigating]. The ability of LLMs to perform zero-shot stance prediction through simple prompting, in particular, dramatically improved the task’s accessibility and performance [18-ding2024edda].
Following this paradigm shift, LLM-centric research has rapidly expanded, moving beyond direct prompting or CoT reasoning toward increasingly sophisticated methodologies [21-zhang2023investigating, 22-zhang2022would]. These include fine-tuning with external knowledge [19-li2023stance], prompt-based multi-perspective reasoning [23-ding2024cross, 17-wang2025more], formalizing logic structures [24-zhang2024knowledge], and designing multi-agent collaborative systems [20-lan2024stance, 25-wang2024deem]. Despite this rapid progress, current approaches continue to treat users as a homogeneous group, overlooking the personal traits and discourse tendencies that shape stance expression. This simplification limits the model’s ability to interpret conflicting viewpoints, capture personalized reactions, and reason about stance in realistic multi-party interactions. Bridging this gap requires moving beyond content-centered modeling toward a genuinely user-centric perspective.
Figure 2: Data construction pipeline of U-MStance.
3 Dataset
This section outlines the construction and key attributes of our U-MStance dataset, comprising 40,003 annotated texts and images. As illustrated in Figure 2, our data construction pipeline ensures high data quality and relevance, involving data collection and data processing to ensure high data quality. Subsequently, data analysis is conducted to profile the dataset’s composition and validate its quality.
3.1 Data Construction
Dataset Collection. To capture realistic user interactions, we sourced conversational data from Reddit [^1] https://www.reddit.com, a large and diverse online forum, using the PRAW library. We began by identifying initial posts related to six high-engagement, real-world targets: Trump, Biden, Tesla, BMW, Costco, and Bitcoin. Posts were selected based on popularity metrics like upvotes and comment counts to ensure a sufficient volume of conversational replies for in-depth annotation. We also collected each user’s historical posts and comments to enrich their personal information.
Filter. The raw data underwent systematic processing to ensure quality and consistency. The process began with a filtering stage that removed content from suspended or deleted user accounts and discarded conversational threads with a depth greater than 9 to maintain relevance.
Pre-annotation. Following this, GPT-4o-mini was employed for automated pre-annotation, assigning a preliminary stance label (Favor, Against, or None) to each text relative to its target.
Human Verification. The automated labels underwent a rigorous human verification process by seven NLP researchers. Annotators were first qualified through two pilot rounds and an expert review. During the main annotation, each instance was independently reviewed by at least two annotators. Disagreements were resolved by a third senior annotator, and the final labels were determined through majority voting. Beyond stance correction, annotators also evaluated the contextual relevance of any associated images. This multi-annotator workflow ensures high label fidelity, incorporates diverse expertise, and reinforces the overall reliability of the dataset.
Table 1: The comparison between our U-MStance with other stance detection datasets. Comment: the number of comment texts; Conv.: support for the conversational context; A-Turns: average turns per conversation; Img: support for the image modality; Target: support for the target type(noun/claim); User: support for user information; -: not mentioned in the original paper.
Dataset Comment Conv. A-Turns Img Target User
SEM16 [2-mohammad2016semeval] 4,870 1.00 noun
P-stance [31-li2021p] 21,574 1.00 noun
COVID-19 [26-glandt2021stance] 6,133 1.00 noun
WT-WT [32-conforti2020will] 51,284 1.00 noun
VAST [3-allaway2020zero] 23,525 1.00 noun
SRQ [4-villa2020stance] 1,348 1.00 noun
CANT-CSD [5-li2023improved] 5,376 11.75 claim
CTSDT [33-li2023contextual] 53,861 - noun
MT-CSD [30-niu2024challenge] 15,876 4.82 noun
C-MTCSD [34-niu2025c] 24,264 3.13 noun
MT2-CSD [35-niu2025mt2] 24,457 4.41 claim
MMVax [37-weinzierl2023identification] 11,300 1.00 noun
MMSD [36-liang2024multi] 17,544 1.00 noun
MmMtCSD [1-niu2024multimodal] 21,340 3.46 claim
U-MStance(Ours) 40,003 4.10 noun
Figure 3: The distribution of stance categories in our U-MStance.
Figure 4: Overflow of our PRISM framework.
3.2 Data Analysis
Data Statistics. Table 1 and Figure 3 provide a statistical overview of U-MStance, contextualizing it among existing stance detection datasets. Our dataset comprises 40,003 annotated instances. Crucially, as highlighted in Table 1, U-MStance is the first dataset in this domain to systematically incorporate user information, offering a foundational resource to address the prevalent issue of user homogeneity. Figure 3 illustrates the distribution of instances across different labels. For model development, the dataset is partitioned into training, validation, and test sets using an 8:1:1 ratio, ensuring a balanced representation for comprehensive evaluation and analysis.
Quality Assessment. To ensure the annotation reliability, we measured inter-annotator agreement using Cohen’s Kappa [38-cohen1960coefficient], achieving an average score of 0.64, indicative of substantial agreement. This consistency was further reinforced through a rigorous post-annotation review, in which senior researchers audited a randomly selected subset of the annotated data to verify accuracy and resolve borderline cases, providing an additional layer of quality assurance.
4 Our PRISM Approach
This section presents a detailed description of our proposed PRISM framework. Given a conversational thread about a global target $T_{g}$, initiated by a post $P=(t_{0},v_{0})$, where $t_{0}$ and $v_{0}$ are the initial text and image. The post is followed by a sequence of comments $C=\{c_{1},...,c_{N}\}$, where each comment $c_{i}$ is a tuple $(u_{i},t_{i},v_{i})$ representing the user, their utterance, and an optional image. Crucially, the input also includes the multimodal historical activity $H_{u_{i}}$ for each user $u_{i}$ involved in the conversation, comprising their past posts and comments. Our objective is to predict the stance label $y_{N}$ of the final comment $c_{N}$ with respect to $T_{g}$.
As illustrated in Figure 4, PRISM comprises three key components: 1) Longitudinal User Persona Distillation infers each user’s Big Five-based personality representation from historical posts and comments; 2) Rationalized Cross-Modal Grounding aligns multimodal cues with conversational context using CoT reasoning via an MLLM to generate intent‑aware captions; 3) Mutual Task Reinforcement jointly optimizes stance detection and stance-aware response generation under a multitask objective.
4.1 Longitudinal User Persona Distillation
Stance is not an isolated behavior but an externalized reflection of a user’s underlying personal tendencies. Thus, each user is modeled as a persistent yet evolving entity whose stance expression is influenced by stable personality traits. For this purpose, we adopt the empirically validated Big Five (OCEAN) personality model [44-jiang2024personallm] from psychology. This model characterizes personality along five fundamental dimensions: Openness, Conscientiousness, Extraversion, Agreeableness, and Neuroticism. To infer these traits for user $u_{N}$, we first aggregate their complete multimodal history, $H_{u_{N}}$, which comprises all of their past posts and comments.
The distillation process leverages the full power of the MLLM $\mathcal{M}$ to reason over long-form, raw multimodal data, allowing it to capture the nuances of the user. To guide the inference, we design a detailed system instruction, $\mathcal{I}_{\text{ocean}}$. This prompt meticulously defines each of the five personality traits and specifies the output format: a numerical score from 1 to 5 for each trait, representing qualitative levels from very low to very high. The MLLM then generates a structured persona representation, $p_{u_{N}}$, composed of five numerical ratings. This process is formally expressed as:
$p_{u_{N}}$ $=\{r^{(O)},r^{(C)},r^{(E)},r^{(A)},r^{(N)}\}_{u_{N}}$ (1)
$=\mathcal{M}(\mathcal{I}_{ocean},H_{u_{N}})$
where each rating $r\in\{1,2,3,4,5\}$. The resulting profile $p_{u_{N}}$ is a concise, structured text string that serves as a crucial, personalized context vector. This profile is explicitly injected into the prompt for the final stance detection task, ensuring the model’s prediction is conditioned on the stable identity of the user in question.
4.2 Rationalized Cross-Modal Grounding
Images in online conversations are not merely decorative; they are communicative acts whose pragmatic meaning is determined by the conversational context. To decipher this context-dependent rhetorical intent, we propose the Rationalized Cross-Modal Grounding (RCMG) module, which generates an intent-aware caption for each image in the conversation. This module employs a two-stage reasoning process, inspired by Chain-of-Thought [45-wei2022chain], to bridge the semantic-pragmatic gap.
Objective Description. The MLLM $\mathcal{M}$ first performs a context-free analysis of the image $v_{i}$, producing a factual, objective description $x_{i}^{obj}$ that grounds reasoning in the literal visual content. This step is guided by a descriptive instruction $I_{desc}$:
$x^{obj}_{i}=\mathcal{M}(I_{desc},v_{i})$ (2)
Intent-aware Interpretation. The objective description $x_{i}^{obj}$ serves as an intermediate reasoning step to support the more complex task of intent inference. At this stage, the MLLM is prompted to infer the speaker’s underlying motivation. We construct a prompt with a specific instruction $\mathcal{I}_{int}$, the image $v_{i}$, its objective description $x_{i}^{obj}$, and the textual content of the current conversation thread $C_{text}$. The model then generates the final intent-aware caption $\hat{x}_{i}$:
$\hat{x}_{i}=\mathcal{M}(I_{int},v_{i},x_{i}^{obj},C_{text})$ (3)
The resulting caption $\hat{x}_{i}$ captures the rhetorical function of the image within the conversational context.
4.3 Mutual Task Reinforcement
To cultivate a deep, user-centric understanding of stance, PRISM adopts a multitask learning paradigm that jointly optimizes two inherently correlated tasks. The central intuition is that by training the model to generate a user’s response (the auxiliary task), it gains a richer understanding of the pragmatic and user-specific cues necessary to accurately predict the stance itself (the primary task). This mutual reinforcement acts as a strong regularizer, enhancing the robustness and prediction accuracy.
Stance Detection (Primary Task). The main objective is to predict the stance label $y_{N}$ of the final comment $c_{N}$. To accomplish this, the MLLM is provided with a comprehensive set of conditioning information including the global target $T_{g}$, the persona representation of the final user $p_{u_{N}}$, the multimodal conversational comment $C$, and the set of all intent-aware captions $\mathcal{X}$ generated by the RCMG module for every image in the conversation. This combination of raw multimodal data and explicit textual reasoning allows the model to integrate these complementary signals. Formally, we denote the full conditioning set as:
$X_{cls}=\{T_{g},p_{u_{N}},C,\mathcal{X}\}$ (4)
The classification loss is the negative log-probability of generating the ground-truth label $y_{N}$.
$\mathcal{L}_{cls}=-\log P(y_{N}|X_{cls})$ (5)
Stance-aware Response Generation (Auxiliary Task). The auxiliary task involves generating the text utterance of the final comment $c_{N}$. This task is conditioned on all preceding context, including the original multimodal conversational comments up to the penultimate turn $C_{<N}=\{c_{1},...,c_{N-1}\}$ and the corresponding set of intent-aware captions $\mathcal{X}_{<N}$ for all images appearing in the preceding turns. The conditioning information also includes the final user’s persona $p_{u_{N}}$ and their ground-truth stance $y_{N}$. We denote this conditioning set as:
$X_{gen}=\{T_{g},y_{N},p_{u_{N}},C_{<N},\mathcal{X}_{<N}\}$ (6)
The generation loss is the standard negative log-likelihood over the tokens $w_{j}$ of the utterance $t_{N}$:
$\mathcal{L}_{gen}=-\sum_{j=1}^{|t_{N}|}\log P(w_{j}|w_{<j},X_{gen})$ (7)
Joint Optimization. The final training objective is a weighted sum of the two losses:
$\mathcal{L}_{total}=\lambda\mathcal{L}_{cls}+(1-\lambda)\mathcal{L}_{gen}$ (8)
This mutual reinforcement ensures PRISM learns a holistic model of stance expression.
Table 2: In-target main results (%), where the methods based on LLMs involve testing through direct questioning. Best and second-best scores are in bold and underline, respectively. indicates our PRISM model significantly outperforms baselines ($p<0.05$). The dashed line separates fine-tuned from non-fine-tuned methods. w/ denotes the backbone. Ag, Fa, Avg represent F1-against, F1-favor, and F1-avg.
Modality Method Trump Biden Tesla BMW Costco Bitcoin Overall
Ag Fa Avg Ag Fa Avg Ag Fa Avg Ag Fa Avg Ag Fa Avg Ag Fa Avg
Text-only BERT 65.94 11.57 38.75 65.31 65.31 54.46 56.07 39.61 47.85 64.22 53.66 58.94 51.83 41.94 46.89 36.96 63.43 50.19 49.51
RoBERTa 66.81 24.19 45.51 64.37 64.37 58.39 57.59 47.62 52.61 64.45 54.07 59.26 62.14 54.94 58.54 49.24 68.89 59.07 55.56
PoliBERTweet 60.00 30.00 45.00 58.70 57.58 58.14 41.67 34.69 38.18 60.00 54.28 57.14 57.56 51.41 54.48 28.07 58.97 43.52 49.41
LLaMA2 24.32 10.26 17.29 23.81 17.50 20.65 32.62 32.50 32.56 35.74 33.45 34.60 30.73 38.76 34.75 26.60 42.74 34.67 29.09
GPT-3.5 62.27 27.59 44.93 51.14 58.80 54.97 63.92 45.18 54.55 60.73 50.22 55.47 62.03 52.31 57.17 47.97 65.44 56.71 53.97
GPT-4 67.52 38.71 53.11 56.28 67.34 61.81 66.34 55.19 60.77 67.76 61.32 64.54 63.35 60.95 62.15 56.44 67.73 62.08 60.74
Multimodal BERT+ViT 63.98 25.85 44.91 49.33 61.45 55.39 56.46 40.27 48.37 65.00 55.56 60.28 62.84 51.90 57.37 34.67 67.37 51.02 52.89
TMPT 66.02 27.91 46.97 50.69 62.07 56.38 53.08 34.97 44.02 61.30 58.38 59.84 60.68 52.39 56.54 44.07 66.31 55.18 53.16
Qwen-VL 62.50 24.33 43.42 49.49 54.63 52.06 50.44 44.23 47.34 53.87 42.64 48.26 62.60 54.82 58.71 50.45 69.60 60.03 51.64
LLaVa 53.18 19.94 36.56 30.23 46.15 38.19 43.09 31.95 37.52 40.30 26.89 33.60 44.92 31.39 38.16 22.88 55.13 39.00 37.17
MiMo 59.08 21.05 40.07 40.63 48.84 44.73 49.53 37.91 43.72 52.60 30.65 41.62 54.76 38.06 46.41 39.02 57.22 48.12 44.11
GPT4-1 69.94 47.74 58.84 55.49 66.48 60.99 74.14 61.37 67.75 70.20 70.52 70.36 74.88 72.40 73.64 58.94 72.82 65.88 66.24
PRISM (ours) 73.84 48.05 60.94 58.82 73.22 66.02 72.06 63.64 67.85 75.00 67.61 71.30 72.44 65.72 69.08 72.00 79.45 75.72 68.49
Backbones - w/ LLaVA 70.47 37.42 53.94 40.00 63.00 51.50 67.16 49.79 58.47 67.63 68.44 68.04 69.93 63.60 66.77 60.91 73.47 67.19 60.99
- w/ MiMo 74.40 53.01 63.71 60.61 69.96 65.28 75.12 61.22 68.17 75.80 76.86 76.33 71.72 68.36 70.04 73.22 81.53 77.38 70.15
5 Experiment
In this section, we present the baseline models evaluated on our proposed dataset, encompassing the experimental setup with implementation details and evaluation metrics, a description of the baseline methods, and a summary of the experimental results achieved by our PRISM framework.
5.1 Experimental Setup
Implement details. Our framework is built upon the open-source Qwen2.5-VL-7B model [48-bai2025qwen2], and is fine-tuned end-to-end for 5 epochs using our multitask objective, with the loss-weighting hyperparameter $\lambda$ set to 0.7. Training is performed with the AdamW optimizer [51-adam2014method], an initial learning rate of $1\times 10^{-5}$, a cosine decay schedule, and a global batch size of 16 achieved through gradient accumulation. All experiments were run on a server with four NVIDIA A100 (40 GB) GPUs using PyTorch and the Hugging Face Transformers library.
Evaluation metrics. We adopt the average F1 score (F1-avg) as the primary evaluation metric, following [1-niu2024multimodal] and [50-niu2025c]. F1-avg is computed as the mean of F1 scores for the against and favor stances, denoted as F1-against and F1-favor. In addition, we also report the overall F1 score for comprehensive comparison.
5.2 Baseline Methods
We conduct extensive experiments with a range of existing methods, categorized into text-only and multimodal approaches. These baselines are evaluated across diverse stance detection scenarios, providing a thorough analysis of their effectiveness and highlighting the benefits of incorporating multimodal information into the stance detection.
Text-only baselines. We evaluate the following methods: (1) Pre-trained BERT [39-devlin2019bert], fine-tuned on the training data; (2) Pre-trained RoBERTa [40-liu2019roberta], an enhancement over BERT with larger batch sizes and more training data; (3) PoliBERTweet [43-kawintiranon2022polibertweet], a BERTweet-base model specialized for Twitter political posts. We also incorporate LLMs, including (4) LLaMA-2-7B (LLaMA2) [49-touvron2023llama], (5) GPT-3.5222https://platform.openai.com/docs/models/gpt-3-5 and (6) GPT-4333https://platform.openai.com/docs/models/gpt-4-turbo. Multimodal baselines. We include: (1) BERT-ViT [36-liang2024multi] which uses BERT for textual encoding and ViT for visual encoding; (2) TMPT [36-liang2024multi] employing targeted prompts for both pre-trained language and visual models; (3) LLaVA-1.5-7B (LLaVA) [47-liu2024improved]; (4) Qwen2.5-VL-7B (Qwen-VL) [48-bai2025qwen2]; (5) MiMo-VL-7B-SFT (MiMo) [46-xiaomi2025mimo]; and (6) GPT4-1444https://platform.openai.com/docs/models/gpt-4.1.
Table 3: Comparison of different models for cross-target stance detection, where the methods based on LLMs involve testing through direct questioning. Best and second-best scores are in bold and underline, respectively. indicates our PRISM model significantly outperforms baselines ($p<0.05$). The dashed line separates fine-tuned from non-fine-tuned methods. w/ denotes the backbone. Ag, Fa, Avg represent F1-against, F1-favor, and F1-avg.
Modality Method Trump Biden Tesla BMW Costco Bitcoin Overall
Ag Fa Avg Ag Fa Avg Ag Fa Avg Ag Fa Avg Ag Fa Avg Ag Fa Avg
Text-only BERT 1.72 19.76 10.74 23.75 25.18 24.27 26.55 16.21 21.38 39.46 36.69 34.02 38.06 16.35 28.99 21.27 29.30 25.28 24.11
RoBERTa 14.69 25.94 20.32 36.72 40.15 38.43 35.19 21.39 28.29 53.04 35.78 47.15 2.89 5.66 6.44 35.08 29.64 32.36 28.83
PoliBERTweet 42.85 3.84 23.35 33.96 43.64 38.80 35.46 30.93 33.19 48.49 31.03 39.76 9.71 20.00 14.85 25.00 36.36 30.68 30.11
Multimodal BERT+ViT 6.83 23.91 15.37 31.21 28.69 29.94 23.58 16.00 19.79 44.50 32.71 38.60 7.58 23.77 15.68 27.11 21.82 28.19 24.60
TMPT 7.90 25.12 16.52 31.75 29.66 30.75 13.98 7.41 10.69 47.39 19.86 33.62 11.63 15.31 13.47 30.61 38.68 34.65 23.28
PRISM (ours) 59.50 28.43 43.97 48.54 61.59 55.06 62.30 56.89 59.60 66.06 51.02 58.54 45.16 57.75 51.45 60.50 67.70 64.10 55.45
Backbones - w/ LLaVA 53.31 22.35 37.83 35.24 59.63 47.43 59.39 49.81 54.60 58.52 58.82 58.67 46.63 43.26 44.94 56.07 61.02 58.55 50.34
- w/ MiMo 60.73 34.39 47.56 43.43 67.03 55.23 67.90 56.37 62.14 64.47 53.63 59.05 47.47 56.92 52.20 60.26 68.90 64.58 56.79
5.3 Main Results
In-Target Stance Detection. We report the experimental results on the U-MStance dataset under the in-target setup, as shown in Table 2, where the training and testing sets share identical targets. The results reveal several key observations. Notably, our proposed PRISM framework achieves the highest overall performance with an F1-avg of 68.49%, which substantially outperforms multimodal GPT4-1 with 66.24% and the text-only GPT-4 with 60.74%. This improvement highlights PRISM’s effectiveness and efficiency in capturing multimodal stance representations even with a relatively lightweight backbone. The results also confirm the importance of visual signals for stance understanding, as evidenced by consistent gains from multimodal models such as BERT-ViT compared to BERT, and GPT4-1 compared to GPT-4. However, social media image often convey subtle or context-dependent meanings that are challenging to interpret from surface-level visual features. By incorporating rationalized cross-modal grounding, PRISM effectively aligns textual and visual semantics, producing more accurate and context-aware stance predictions. Interestingly, the large models in our experiments generally underperform compared to smaller, specialized models. This may be because, despite their strong general language understanding, they are less effective at capturing fine-grained, conversation-specific, and multimodal cues that are critical for stance detection. In contrast, our PRISM leverages an LLM as its backbone, but through the integration of user persona and multimodality modeling, it effectively captures these nuanced cues, resulting in superior performance.
Cross-Target Stance Detection. We further evaluate generalization under the cross-target setting, where targets in the test set are unseen during training, as summarized in Table 3. Specifically, we conduct cross-target evaluations between related pairs such as Trump-Biden, Tesla-BMW, and Costco-Bitcoin, training on one and testing on the other. This scenario is more challenging, requiring the model to transfer stance understanding across targets with differing semantics and discourse style. LLMs are evaluated directly without additional fine-tuning, yielding performance identical to the in-target setup. We can observe that most smaller models suffer a notable degradation in performance, revealing their limited capacity to generalize beyond the specific targets encountered during training. Remarkably, our proposed PRISM framework maintains a relatively high and stable performance across all cross-target evaluations, demonstrating superior robustness and adaptability. This advantage stems from two components: the integration of user persona modeling, which allows the model to capture consistent user-level stance tendencies across topics, and the mutual task reinforcement mechanism, which jointly optimizes stance prediction and rationalized cross-modal grounding. Together, these components enable PRISM to establish a more target-invariant representation of stance, thereby enhancing its generalization capability across heterogeneous contexts.
5.4 Impact of Different Backbone
We further assess the adaptability of PRISM by integrating it with different MLLM backbones under both the in-target and cross-target settings, as illustrated in Table 2 and Table 3. Specifically, we replace the original Qwen-VL backbone with LLaVA and MiMo to examine how backbone choice affects performance. Across both evaluation setups, MiMo provides consistent performance gains across both evaluation setup. This finding demonstrates not only the effectiveness but also the generalizability of our framework, showing that PRISM can seamlessly adapt to stronger multimodal backbones while preserving its core advantages. The gains from MiMo likely stem from its enhanced multimodal reasoning capabilities and more robust cross-domain representation learning. These properties align well with PRISM’s design, enabling more precise modeling of nuanced stance cues and improving the transferability of stance understanding across heterogeneous targets.
Table 4: Experimental results (%) of the ablation study on in-target and cross-target multimodal conversational stance detection. w/o means without.
Trump Biden Tesla BMW Costco Bitcoin
In-Target
PRISM 60.94 66.02 67.85 71.30 69.08 75.72
-w/o Persona 55.75 64.18 67.50 69.27 66.74 73.69
-w/o Intent 50.81 57.30 56.45 65.34 61.60 72.59
-w/o Mutual 59.64 64.34 67.07 68.97 68.89 73.34
Cross-Target
PRISM 43.97 55.06 59.60 58.54 51.45 64.10
-w/o Persona 40.82 53.34 58.34 55.17 49.72 61.51
-w/o Intent 40.99 45.19 41.90 48.30 47.84 53.66
-w/o Mutual 43.63 52.16 58.80 54.01 49.88 63.91
Figure 5: Case study comparing PRISM with baseline models.
5.5 Ablation Study
To evaluate the contribution of each component in our PRISM framework, we conduct a comprehensive ablation study with the results presented in Table 4. Removing the persona representation $p_{u_{N}}$ (w/o Persona) yields a consistent performance drop of approximately 2%-3%, underscoring the significance of modeling user-specific characteristics. This component enables the framework to capture individualized traits, thereby enhancing its ability to accurately interpret user-centric stance expressions across conversational contexts. Eliminating the intent-aware caption $\hat{x}$ (w/o Intent) results in a more substantial degradation, highlighting the crucial role of the rationalized cross-modal grounding module. By explicitly aligning the pragmatic intent of images with the accompanying textual context, this component enables the model to reason about subtle or implicit visual cues, which are often critical for correctly inferring stance in multimodal conversations. Lastly, discarding the mutual task reinforcement mechanism and reverting to a single-task stance detection objective (w/o Mutual) also leads to a noticeable decline in performance. This confirms that jointly optimizing stance detection with auxiliary generation task encourages beneficial interactions between multimodal understanding and stance reasoning, ultimately producing more robust and transferable representations.
5.6 Case Study
To visually demonstrate the superiority of our method and the inherent challenges of the task, we conduct a case study, as illustrated in the figure 5. As shown in Case (a), both GPT4-1 and LLaVA fail to detect the sarcastic implication conveyed jointly by the text and image, incorrectly assigning a neutral stance. These models tend to rely on surface-level cues and struggle with recognizing pragmatic intent. In contrast, our PRISM accurately predicts an Against stance. This improvement stems from its rationalized cross-modal grounding, which captures the user’s implicit criticism expressed through the comparison embedded in the image, as well as its persona-aware interpretation that contextualizes the user’s communicative tendencies.
Meanwhile, Case (b) presents an even more subtle scenario, where sarcasm is intertwined with the user’s disposition. Again, GPT4-1 and LLaVA misclassify the stance as neutral, likely due to their lack of user-level modeling and reliance solely on the current conversational snippet. PRISM, however, leverages its persona module to interpret the user’s longitudinal behavioral patterns. As shown in the PRISM persona visualization, the user exhibits a prominently high Neuroticism score, reflecting a more emotionally expressive and critical communication style. Incorporating such personality cues enables PRISM to infer the user’s consistent critical orientation, allowing it to correctly identify the Against stance even in contexts where the sarcastic intent is subtle and easily overlooked.
Figure 6: Performance of different depths. Conversational depth: S (1-2 turns), M (3-4 turns), L ($\geq$5 turns).
5.7 Impact of Conversation Depth
We further investigate how conversational complexity influences stance detection by evaluating PRISM across varying conversation depths, as shown in Figure 6. Overall, performance shows a gradual decline as depth increases, which is expected since deeper conversations involve more intricate discourse structures, longer-range dependencies, and greater contextual ambiguity, making the final stance harder to infer. Notably, for the targets Costco and Bitcoin, PRISM maintains relatively stable performance across depths. This robustness is largely attributed to the rationalized cross-modal grounding module, whose intent-aware image captions help align visual and textual cues and preserve salient multimodal information, thereby mitigating the challenges posed by longer conversational contexts.
6 Conclusion
This paper presents two critical limitations in MCSD, including pseudo-multimodality, where visual cues are limited to source posts, and user homogeneity, which neglects individual differences. To address these, we propose U-MStance, the first user-centric MCSD dataset with multimodal and user information. Furthermore, we design PRISM, a framework that distills stable user personas, grounds image intent, and reinforces stance understanding through multitask learning. Entensive Experiments on U-MStance demonstrate that PRISM consistently outperforms strong baselines and exhibits robust generalization.

### v2

@@
 Finally, a mutual task reinforcement mechanism is employed to jointly optimize stance detection and stance-aware response generation for bidirectional knowledge transfer.
-Experiments on U‑MStance demonstrate that PRISM yields significant gains over strong baselines, underscoring the effectiveness of user-centric and context-grounded multimodal reasoning for realistic stance understanding.
+Experiments on U-MStance demonstrate that PRISM yields significant gains over strong baselines, underscoring the effectiveness of user-centric and context-grounded multimodal reasoning for realistic stance understanding.
 1 Introduction
-Social media have become a dominant platform for the expression of opinions, making their analysis crucial to understand divisive societal topics such as politics, public health, and social justice [26-glandt2021stance, 27-liang2022zero].
+Social media have become a dominant platform for the expression of opinions, making their analysis crucial to understand divisive societal topics such as politics, public health, and social justice [13, 25].
 Stance detection, the task of identifying a user’s position toward a target, has become vital for charting public discourse and polarization.
-As online interactions grow more complex, the field has evolved from simple textual analysis to modeling rich, multi-party conversational contexts [28-li2023improved, 29-li2023contextual, 30-niu2024challenge].
+As online interactions grow more complex, the field has evolved from simple textual analysis to modeling rich, multi-party conversational contexts [21, 23, 33].
 Figure 1: Comparison of a conventional approach for MCSD (left) with our proposed PRISM framework (right).
-The recent shift towards Multimodal Conversational Stance Detection (MCSD), exemplified by the MmMtCSD dataset [1-niu2024multimodal], marks a crucial step forward.
+The recent shift towards Multimodal Conversational Stance Detection (MCSD), exemplified by the MmMtCSD dataset [31], marks a crucial step forward.
 However, despite this progress, two fundamental limitations persist across the field, hindering authentic stance modeling.
@@
 Building upon this critical new dataset, we further propose PRISM, a Persona-Reasoned multImodal Stance Model for MCSD.
-Specifically, PRISM first explores the underlying personal tendencies behind stance expressions and derives longitudinal user persona traits inspired by Big Five (OCEAN) theory [44-jiang2024personallm] from historical posts and comments.
+Specifically, PRISM first explores the underlying personal tendencies behind stance expressions and derives longitudinal user persona traits inspired by Big Five (OCEAN) theory [14] from historical posts and comments.
 Then, it introduces a rationalized cross-modal grounding module, leveraging Chain-of-Thought (CoT) reasoning to explicitly bridge the profound semantic and pragmatic gap between modalities in conversational contexts.
@@
 The development of stance detection datasets has evolved from analyzing isolated texts to modeling complex interactions.
-Early benchmarks like SemEval-2016 Task 6 [2-mohammad2016semeval] and VAST [3-allaway2020zero] focused on sentence-level textual stance detection, establishing foundational models but overlooking the context of dialogue.
-This limitation spurred the creation of conversational datasets such as SRQ [4-villa2020stance] and CANT-CSD [5-li2023improved], which treated stance as a dynamic process unfolding within a conversation.
+Early benchmarks like SemEval-2016 Task 6 [30] and VAST [3] focused on sentence-level textual stance detection, establishing foundational models but overlooking the context of dialogue.
+This limitation spurred the creation of conversational datasets such as SRQ [36] and CANT-CSD [22], which treated stance as a dynamic process unfolding within a conversation.
 The most recent advancement is the shift toward multimodal stance detection, which incorporates visual elements alongside text to better reflect modern social media discourse.
-Despite these advances, even the latest datasets for MCSD, such as MmMtCSD [1-niu2024multimodal], still face two key shortcomings that hinder realistic stance modeling.
-First, multimodality is handled only superficially, with source posts treated as multimodal but conversational replies reduced to text-only.
+However, even the latest datasets for MCSD, such as MmMtCSD [31], still face two key shortcomings that hinder realistic stance modeling.
+First, multimodality is constrained to source posts as static background context, while conversational replies are restricted to text-only.
 In addition, information about users, including their posting history, is largely ignored, overlooking contextual factors that shape an individual’s expressed stance and leaving a substantial gap in building comprehensive and authentic stance detection models.
 2.2 Stance Detection Methods
-Methodological advancements in stance detection has spanned multiple task settings in-target [6-li2021multi], cross-target [7-ding2024cross, 8-liang2021target, 9-wei2019modeling], and zero-shot scenarios [10-allaway2020zero, 11-liang2022zero, 12-zhu2022enhancing].
-Early deep learning approaches primarily fine-tuned pre-trained models such as BERT [13-cambria2018senticnet].
-However, the field was revolutionized by the emergence of Large Language Models (LLMs), whose powerful semantic understanding and adaptability introduced new capabilities [14-cheng2024emotion, 15-cheng2024mips, 16-li2024mitigating].
-The ability of LLMs to perform zero-shot stance prediction through simple prompting, in particular, dramatically improved the task’s accessibility and performance [18-ding2024edda].
-Following this paradigm shift, LLM-centric research has rapidly expanded, moving beyond direct prompting or CoT reasoning toward increasingly sophisticated methodologies [21-zhang2023investigating, 22-zhang2022would].
-These include fine-tuning with external knowledge [19-li2023stance], prompt-based multi-perspective reasoning [23-ding2024cross, 17-wang2025more], formalizing logic structures [24-zhang2024knowledge], and designing multi-agent collaborative systems [20-lan2024stance, 25-wang2024deem].
-Despite this rapid progress, current approaches continue to treat users as a homogeneous group, overlooking the personal traits and discourse tendencies that shape stance expression.
-This simplification limits the model’s ability to interpret conflicting viewpoints, capture personalized reactions, and reason about stance in realistic multi-party interactions.
-Bridging this gap requires moving beyond content-centered modeling toward a genuinely user-centric perspective.
+Methodological advancements in stance detection has spanned multiple task settings in-target [19], cross-target [11, 26, 40], and zero-shot scenarios [2, 24, 46].
+Early deep learning approaches primarily fine-tuned pre-trained models such as BERT [5].
+However, the field was revolutionized by the emergence of Large Language Models (LLMs), whose powerful semantic understanding and adaptability introduced new capabilities [6, 7, 18, 12].
+Following this paradigm shift, LLM-centric research has moved beyond direct prompting or CoT reasoning toward increasingly sophisticated methodologies [45, 44].
+These include fine-tuning with external knowledge [17], prompt-based multi-perspective reasoning [11, 37], formalizing logic structures [43], and designing multi-agent collaborative systems [16, 38].
+Despite this rapid progress, current approaches often treat users as homogeneous, overlooking the personal traits and discourse tendencies that shape stance expression.
+This limits the interpretation of personalized reactions and conflicting views, necessitating a shift from content-centered to user-centric modeling.
 Figure 2: Data construction pipeline of U-MStance.
 3 Dataset
-This section outlines the construction and key attributes of our U-MStance dataset, comprising 40,003 annotated texts and images.
-As illustrated in Figure 2, our data construction pipeline ensures high data quality and relevance, involving data collection and data processing to ensure high data quality.
+This section outlines the construction and key attributes of our U-MStance dataset.
+As illustrated in Figure 2, our data construction pipeline includes collection, processing and verification to ensure data quality and relevance.
 Subsequently, data analysis is conducted to profile the dataset’s composition and validate its quality.
@@
 To capture realistic user interactions, we sourced conversational data from Reddit [^1] https://www.reddit.com, a large and diverse online forum, using the PRAW library.
-We began by identifying initial posts related to six high-engagement, real-world targets: Trump, Biden, Tesla, BMW, Costco, and Bitcoin.
-Posts were selected based on popularity metrics like upvotes and comment counts to ensure a sufficient volume of conversational replies for in-depth annotation.
+Covering the period from 2017 to 2025, we initially identified 835 posts related to six high-engagement, real-world targets: Trump, Biden, Tesla, BMW, Costco, and Bitcoin.
+These posts were selected based on popularity metrics like upvotes and comment counts to ensure a sufficient volume of conversational replies.
 We also collected each user’s historical posts and comments to enrich their personal information.
@@
 The process began with a filtering stage that removed content from suspended or deleted user accounts and discarded conversational threads with a depth greater than 9 to maintain relevance.
-Pre-annotation. Following this, GPT-4o-mini was employed for automated pre-annotation, assigning a preliminary stance label (Favor, Against, or None) to each text relative to its target.
+Pre-annotation. Following this, GPT-4o-mini was employed for automated pre-annotation, assigning a preliminary stance label (Favor, Against, or None) to each instance relative to its target.
 Human Verification.
@@
 Beyond stance correction, annotators also evaluated the contextual relevance of any associated images.
-This multi-annotator workflow ensures high label fidelity, incorporates diverse expertise, and reinforces the overall reliability of the dataset.
-Table 1: The comparison between our U-MStance with other stance detection datasets.
-Comment: the number of comment texts; Conv.: support for the conversational context; A-Turns: average turns per conversation; Img: support for the image modality; Target: support for the target type(noun/claim); User: support for user information; -: not mentioned in the original paper.
-Dataset Comment Conv. A-Turns Img Target User
-SEM16 [2-mohammad2016semeval] 4,870 1.00 noun
-P-stance [31-li2021p] 21,574 1.00 noun
-COVID-19 [26-glandt2021stance] 6,133 1.00 noun
-WT-WT [32-conforti2020will] 51,284 1.00 noun
-VAST [3-allaway2020zero] 23,525 1.00 noun
-SRQ [4-villa2020stance] 1,348 1.00 noun
-CANT-CSD [5-li2023improved] 5,376 11.75 claim
-CTSDT [33-li2023contextual] 53,861 - noun
-MT-CSD [30-niu2024challenge] 15,876 4.82 noun
-C-MTCSD [34-niu2025c] 24,264 3.13 noun
-MT2-CSD [35-niu2025mt2] 24,457 4.41 claim
-MMVax [37-weinzierl2023identification] 11,300 1.00 noun
-MMSD [36-liang2024multi] 17,544 1.00 noun
-MmMtCSD [1-niu2024multimodal] 21,340 3.46 claim
-U-MStance(Ours) 40,003 4.10 noun
-Figure 3: The distribution of stance categories in our U-MStance.
-Figure 4: Overflow of our PRISM framework.
+This multi-annotator workflow ensures high label fidelity, and reinforces the overall reliability of the dataset.
+Table 1: Comparison of U-MStance with existing datasets.
+Size: the number of annotated instances; Target: the number of the targets; Conv.: support for the conversational context; Vision-Related: percentage of vision-integrated instances; User: support for user information; -: not applicable.
+Dataset Size Target Conv. Vision-Related User
+SEM16 [30] 4,870 6
+P-stance [20] 21,574 3
+COVID-19 [13] 6,133 4
+WT-WT [9] 51,284 5
+VAST [3] 23,525 -
+SRQ [36] 1,348 4
+CANT-CSD [22] 5,376 1
+CTSDT [23] 53,861 1
+MT-CSD [33] 15,876 5
+C-MTCSD [34] 24,264 5
+MT2-CSD [32] 24,457 6
+MMVax [41] 11,300 - 100%
+MMSD [27] 17,544 5 100%
+MmMtCSD [31] 21,340 3 65.99%
+U-MStance(Ours) 40,003 6 100%
+Figure 3: Overview of our PRISM framework.
 3.2 Data Analysis
 Data Statistics.
-Table 1 and Figure 3 provide a statistical overview of U-MStance, contextualizing it among existing stance detection datasets.
-Our dataset comprises 40,003 annotated instances.
-Crucially, as highlighted in Table 1, U-MStance is the first dataset in this domain to systematically incorporate user information, offering a foundational resource to address the prevalent issue of user homogeneity.
-Figure 3 illustrates the distribution of instances across different labels.
-For model development, the dataset is partitioned into training, validation, and test sets using an 8:1:1 ratio, ensuring a balanced representation for comprehensive evaluation and analysis.
+Our dataset comprises 40,003 annotated instances and 24952 user profiles.
+Each instance consists of a complete conversational thread accompanied by its associated images within the conversation.
+Crucially, Table 1 provide a comparison between U-MStance and existing datasets.
+U-MStance is the first dataset in this domain to systematically incorporate user information, offering a foundational resource to address the prevalent issue of user homogeneity.
+More comprehensive statistics, including the distributions of labels and conversational depths, are detailed in Appendix A.
+For model development, the dataset is partitioned into training, validation, and test sets using an 8:1:1 ratio.
+We strictly partition the data by cross-referencing timestamps to ensure that the three sets are mutually independent.
 Quality Assessment.
-To ensure the annotation reliability, we measured inter-annotator agreement using Cohen’s Kappa [38-cohen1960coefficient], achieving an average score of 0.64, indicative of substantial agreement.
+To ensure the annotation reliability, we measured inter-annotator agreement using Cohen’s Kappa [8], achieving an average score of 0.64, indicative of substantial agreement.
 This consistency was further reinforced through a rigorous post-annotation review, in which senior researchers audited a randomly selected subset of the annotated data to verify accuracy and resolve borderline cases, providing an additional layer of quality assurance.
@@
 4.1 Longitudinal User Persona Distillation
-Stance is not an isolated behavior but an externalized reflection of a user’s underlying personal tendencies.
-Thus, each user is modeled as a persistent yet evolving entity whose stance expression is influenced by stable personality traits.
-For this purpose, we adopt the empirically validated Big Five (OCEAN) personality model [44-jiang2024personallm] from psychology.
-This model characterizes personality along five fundamental dimensions: Openness, Conscientiousness, Extraversion, Agreeableness, and Neuroticism.
-To infer these traits for user $u_{N}$, we first aggregate their complete multimodal history, $H_{u_{N}}$, which comprises all of their past posts and comments.
+Stance is an externalized reflection of a user’s underlying personal tendencies.
+Thus, each user is modeled as an entity whose stance expression is shaped by stable personality traits.
+For this purpose, we adopt the empirically validated Big Five (OCEAN) personality model [14], which characterizes personality along five fundamental dimensions: Openness, Conscientiousness, Extraversion, Agreeableness, and Neuroticism.
+To infer these traits for user $u_{N}$, we first aggregate their complete history $H_{u_{N}}$, which comprises all historical posts and comments.
 The distillation process leverages the full power of the MLLM $\mathcal{M}$ to reason over long-form, raw multimodal data, allowing it to capture the nuances of the user.
@@
 This profile is explicitly injected into the prompt for the final stance detection task, ensuring the model’s prediction is conditioned on the stable identity of the user in question.
+The prompt template can be found in Table 7 of Appendix C.
 4.2 Rationalized Cross-Modal Grounding
-Images in online conversations are not merely decorative; they are communicative acts whose pragmatic meaning is determined by the conversational context.
-To decipher this context-dependent rhetorical intent, we propose the Rationalized Cross-Modal Grounding (RCMG) module, which generates an intent-aware caption for each image in the conversation.
-This module employs a two-stage reasoning process, inspired by Chain-of-Thought [45-wei2022chain], to bridge the semantic-pragmatic gap.
+Images in online conversations are not decorations but communicative acts whose pragmatic meaning is shaped by the conversational context.
+To decipher this context-dependent rhetorical intent, we propose the Rationalized Cross-Modal Grounding (RCMG) module to generate an intent-aware caption for each image in the conversation.
+Inspired by Chain-of-Thought [39], this module employs a two-stage reasoning process to bridge the semantic-pragmatic gap.
+The complete prompt template can be found in Table 8 of Appendix C.
 Objective Description.
@@
 $\hat{x}_{i}=\mathcal{M}(I_{int},v_{i},x_{i}^{obj},C_{text})$ (3)
-The resulting caption $\hat{x}_{i}$ captures the rhetorical function of the image within the conversational context.
+The resulting caption $\hat{x}_{i}$ captures the rhetorical function of the image within the conversation.
 4.3 Mutual Task Reinforcement
-To cultivate a deep, user-centric understanding of stance, PRISM adopts a multitask learning paradigm that jointly optimizes two inherently correlated tasks.
-The central intuition is that by training the model to generate a user’s response (the auxiliary task), it gains a richer understanding of the pragmatic and user-specific cues necessary to accurately predict the stance itself (the primary task).
-This mutual reinforcement acts as a strong regularizer, enhancing the robustness and prediction accuracy.
+To cultivate a comprehensive, user-centric understanding of stance, PRISM employs a multitask learning paradigm that jointly optimizes two inherently correlated tasks.
+The core intuition is that by training the model to generate a user’s response as an auxiliary task, it derives a deeper understanding of the pragmatic and user-specific cues essential for the primary task of stance prediction.
+This mutual reinforcement serves as a robust regularizer, enhancing the stability and prediction accuracy.
 Stance Detection (Primary Task).
@@
 $X_{cls}=\{T_{g},p_{u_{N}},C,\mathcal{X}\}$ (4)
-The classification loss is the negative log-probability of generating the ground-truth label $y_{N}$.
+The classification loss is the negative log-probability of the ground-truth label $y_{N}$.
 $\mathcal{L}_{cls}=-\log P(y_{N}|X_{cls})$ (5)
@@
 5 Experiment
-In this section, we present the baseline models evaluated on our proposed dataset, encompassing the experimental setup with implementation details and evaluation metrics, a description of the baseline methods, and a summary of the experimental results achieved by our PRISM framework.
 5.1 Experimental Setup
-Implement details.
-Our framework is built upon the open-source Qwen2.5-VL-7B model [48-bai2025qwen2], and is fine-tuned end-to-end for 5 epochs using our multitask objective, with the loss-weighting hyperparameter $\lambda$ set to 0.7.
-Training is performed with the AdamW optimizer [51-adam2014method], an initial learning rate of $1\times 10^{-5}$, a cosine decay schedule, and a global batch size of 16 achieved through gradient accumulation.
+Implementation details.
+Our framework is built on Qwen2.5-VL-7B [4] and fine-tuned for 5 epochs using our multitask objective, with the loss-weighting hyperparameter $\lambda$ set to 0.7.
+We use the AdamW optimizer [1], with an initial learning rate of $1\times 10^{-5}$, a cosine decay schedule, and a global batch size of 16.
 All experiments were run on a server with four NVIDIA A100 (40 GB) GPUs using PyTorch and the Hugging Face Transformers library.
 Evaluation metrics.
-We adopt the average F1 score (F1-avg) as the primary evaluation metric, following [1-niu2024multimodal] and [50-niu2025c].
+We adopt the average F1 score (F1-avg) as the primary evaluation metric, following [31] and [34].
 F1-avg is computed as the mean of F1 scores for the against and favor stances, denoted as F1-against and F1-favor.
-In addition, we also report the overall F1 score for comprehensive comparison.
 5.2 Baseline Methods
 We conduct extensive experiments with a range of existing methods, categorized into text-only and multimodal approaches.
-These baselines are evaluated across diverse stance detection scenarios, providing a thorough analysis of their effectiveness and highlighting the benefits of incorporating multimodal information into the stance detection.
+All baseline models are fine-tuned on our specialized training data, with the exception of the GPT series, which is evaluated in a zero-shot setting.
 Text-only baselines.
-We evaluate the following methods: (1) Pre-trained BERT [39-devlin2019bert], fine-tuned on the training data; (2) Pre-trained RoBERTa [40-liu2019roberta], an enhancement over BERT with larger batch sizes and more training data; (3) PoliBERTweet [43-kawintiranon2022polibertweet], a BERTweet-base model specialized for Twitter political posts.
-We also incorporate LLMs, including (4) LLaMA-2-7B (LLaMA2) [49-touvron2023llama], (5) GPT-3.5222https://platform.openai.com/docs/models/gpt-3-5 and (6) GPT-4333https://platform.openai.com/docs/models/gpt-4-turbo.
+We evaluate the following methods: BERT [10], RoBERTa [29], PoliBERTweet [15], LLaMA2 [35], GPT-3.5 and GPT-4.
 Multimodal baselines.
-We include: (1) BERT-ViT [36-liang2024multi] which uses BERT for textual encoding and ViT for visual encoding; (2) TMPT [36-liang2024multi] employing targeted prompts for both pre-trained language and visual models; (3) LLaVA-1.5-7B (LLaVA) [47-liu2024improved]; (4) Qwen2.5-VL-7B (Qwen-VL) [48-bai2025qwen2]; (5) MiMo-VL-7B-SFT (MiMo) [46-xiaomi2025mimo]; and (6) GPT4-1444https://platform.openai.com/docs/models/gpt-4.1.
+We include: BERT-ViT [27], TMPT [27], LLaVA [28], Qwen-VL [4], MiMo [42] and GPT4-1.
+For more details, please refer to Appendix B.
 Table 3: Comparison of different models for cross-target stance detection, where the methods based on LLMs involve testing through direct questioning.
@@
 In-Target Stance Detection.
-We report the experimental results on the U-MStance dataset under the in-target setup, as shown in Table 2, where the training and testing sets share identical targets.
-The results reveal several key observations.
-Notably, our proposed PRISM framework achieves the highest overall performance with an F1-avg of 68.49%, which substantially outperforms multimodal GPT4-1 with 66.24% and the text-only GPT-4 with 60.74%.
+We report the experimental results on the U-MStance dataset under the in-target setup, where the training and testing sets share identical targets.
+As shown in Table 2, our proposed PRISM framework achieves the highest overall performance with an F1-avg of 68.49%, which substantially outperforms multimodal GPT4-1 with 66.24% and the text-only GPT-4 with 60.74%.
 This improvement highlights PRISM’s effectiveness and efficiency in capturing multimodal stance representations even with a relatively lightweight backbone.
 The results also confirm the importance of visual signals for stance understanding, as evidenced by consistent gains from multimodal models such as BERT-ViT compared to BERT, and GPT4-1 compared to GPT-4.
-However, social media image often convey subtle or context-dependent meanings that are challenging to interpret from surface-level visual features.
 By incorporating rationalized cross-modal grounding, PRISM effectively aligns textual and visual semantics, producing more accurate and context-aware stance predictions.
-Interestingly, the large models in our experiments generally underperform compared to smaller, specialized models.
-This may be because, despite their strong general language understanding, they are less effective at capturing fine-grained, conversation-specific, and multimodal cues that are critical for stance detection.
-In contrast, our PRISM leverages an LLM as its backbone, but through the integration of user persona and multimodality modeling, it effectively captures these nuanced cues, resulting in superior performance.
+Interestingly, the large models, such as LLAVA and MiMo, generally underperform compared to smaller, specialized models like TMPT.
+Despite their strong general language understanding, these larger models may be less effective at capturing fine-grained, conversation-specific, and multimodal cues that are critical for stance detection.
+In contrast, our PRISM leverages user persona and multimodal modeling to effectively capture nuanced cues, enabling its LLM backbone to achieve superior performance.
 Cross-Target Stance Detection.
 We further evaluate generalization under the cross-target setting, where targets in the test set are unseen during training, as summarized in Table 3.
-Specifically, we conduct cross-target evaluations between related pairs such as Trump-Biden, Tesla-BMW, and Costco-Bitcoin, training on one and testing on the other.
+Specifically, we conduct cross-target evaluations between pairs such as Trump-Biden, Tesla-BMW, and Costco-Bitcoin, training on one and testing on the other.
 This scenario is more challenging, requiring the model to transfer stance understanding across targets with differing semantics and discourse style.
-LLMs are evaluated directly without additional fine-tuning, yielding performance identical to the in-target setup.
 We can observe that most smaller models suffer a notable degradation in performance, revealing their limited capacity to generalize beyond the specific targets encountered during training.
@@
 Specifically, we replace the original Qwen-VL backbone with LLaVA and MiMo to examine how backbone choice affects performance.
-Across both evaluation setups, MiMo provides consistent performance gains across both evaluation setup.
+Across both evaluation setups, MiMo provides consistent performance gains across both evaluation setups.
 This finding demonstrates not only the effectiveness but also the generalizability of our framework, showing that PRISM can seamlessly adapt to stronger multimodal backbones while preserving its core advantages.
@@
 These properties align well with PRISM’s design, enabling more precise modeling of nuanced stance cues and improving the transferability of stance understanding across heterogeneous targets.
-Table 4: Experimental results (%) of the ablation study on in-target and cross-target multimodal conversational stance detection. w/o means without.
+Table 4: Experimental results (%) of the ablation study. w/o means without.
 Trump Biden Tesla BMW Costco Bitcoin
@@
 PRISM 60.94 66.02 67.85 71.30 69.08 75.72
--w/o Persona 55.75 64.18 67.50 69.27 66.74 73.69
+-w/o Persona 49.56 60.38 61.67 70.02 64.85 72.92
 -w/o Intent 50.81 57.30 56.45 65.34 61.60 72.59
@@
 PRISM 43.97 55.06 59.60 58.54 51.45 64.10
--w/o Persona 40.82 53.34 58.34 55.17 49.72 61.51
+-w/o Persona 40.96 47.16 56.42 55.95 50.13 60.73
 -w/o Intent 40.99 45.19 41.90 48.30 47.84 53.66
@@
 To evaluate the contribution of each component in our PRISM framework, we conduct a comprehensive ablation study with the results presented in Table 4.
-Removing the persona representation $p_{u_{N}}$ (w/o Persona) yields a consistent performance drop of approximately 2%-3%, underscoring the significance of modeling user-specific characteristics.
-This component enables the framework to capture individualized traits, thereby enhancing its ability to accurately interpret user-centric stance expressions across conversational contexts.
-Eliminating the intent-aware caption $\hat{x}$ (w/o Intent) results in a more substantial degradation, highlighting the crucial role of the rationalized cross-modal grounding module.
-By explicitly aligning the pragmatic intent of images with the accompanying textual context, this component enables the model to reason about subtle or implicit visual cues, which are often critical for correctly inferring stance in multimodal conversations.
+Replacing the persona representation $p_{u_{N}}$ with the user’s original raw information (w/o Persona) consistently drops performance across all targets.
+This underscores the critical role of persona distillation in filtering inherent noise and extracting task-relevant characteristics, enabling PRISM to more precisely interpret individualized, user-centric stance expressions.
+Substituting intent-aware captions $\hat{x}$ with raw visual features (w/o Intent) results in a more substantial degradation, highlighting the crucial role of rationalized cross-modal grounding.
+By explicitly aligning the image pragmatic intent with textual context, this component enables the PRISM to correctly infer subtle or implicit visual cues essential for multimodal stance detection.
 Lastly, discarding the mutual task reinforcement mechanism and reverting to a single-task stance detection objective (w/o Mutual) also leads to a noticeable decline in performance.
@@
 5.6 Case Study
-To visually demonstrate the superiority of our method and the inherent challenges of the task, we conduct a case study, as illustrated in the figure 5.
-As shown in Case (a), both GPT4-1 and LLaVA fail to detect the sarcastic implication conveyed jointly by the text and image, incorrectly assigning a neutral stance.
-These models tend to rely on surface-level cues and struggle with recognizing pragmatic intent.
+To visually demonstrate the superiority of our method and the inherent challenges of the task, Figure 4 provides a case study.
+In Case (a), both GPT4-1 and LLaVA fail to detect the sarcastic implication conveyed by joint multimodal cues, incorrectly assigning a neutral stance due to their reliance on surface-level information.
 In contrast, our PRISM accurately predicts an Against stance.
-This improvement stems from its rationalized cross-modal grounding, which captures the user’s implicit criticism expressed through the comparison embedded in the image, as well as its persona-aware interpretation that contextualizes the user’s communicative tendencies.
-Meanwhile, Case (b) presents an even more subtle scenario, where sarcasm is intertwined with the user’s disposition.
-Again, GPT4-1 and LLaVA misclassify the stance as neutral, likely due to their lack of user-level modeling and reliance solely on the current conversational snippet.
-PRISM, however, leverages its persona module to interpret the user’s longitudinal behavioral patterns.
-As shown in the PRISM persona visualization, the user exhibits a prominently high Neuroticism score, reflecting a more emotionally expressive and critical communication style.
-Incorporating such personality cues enables PRISM to infer the user’s consistent critical orientation, allowing it to correctly identify the Against stance even in contexts where the sarcastic intent is subtle and easily overlooked.
+This improvement stems from its rationalized cross-modal grounding, which captures the user’s implicit criticism expressed through the comparison embedded in the image, as well as its persona-aware interpretation of the user’s communicative tendencies.
+Meanwhile, Case (b) presents a more subtle scenario, where sarcasm is intertwined with the user’s disposition.
+Again, GPT4-1 and LLaVA misclassify the stance, likely due to their lack of user-level modeling and reliance solely on the current conversational snippet.
+PRISM, however, correctly identifies the Against stance by leveraging its persona module to interpret longitudinal behavioral patterns.
+Specifically, the persona visualization highlights a high Neuroticism score, reflecting an emotionally expressive and critical style.
+Incorporating such personality cues enables PRISM to infer the user’s consistent critical orientation even when sarcasm is subtle.
 Figure 5: Performance of different depths. Conversational depth: S (1-2 turns), M (3-4 turns), L ($\geq$5 turns).
 5.7 Impact of Conversation Depth
-We further investigate how conversational complexity influences stance detection by evaluating PRISM across varying conversation depths, as shown in Figure 6.
-Overall, performance shows a gradual decline as depth increases, which is expected since deeper conversations involve more intricate discourse structures, longer-range dependencies, and greater contextual ambiguity, making the final stance harder to infer.
+We further investigate evaluate how PRISM’s stance detection performance is influenced by varying conversational depths.
+As shown in Figure 5, performance shows a gradual decline as depth increases, which is expected since deeper conversations involve more intricate discourse structures, longer-range dependencies, and greater contextual ambiguity, making the final stance harder to infer.
 Notably, for the targets Costco and Bitcoin, PRISM maintains relatively stable performance across depths.
@@
 Furthermore, we design PRISM, a framework that distills stable user personas, grounds image intent, and reinforces stance understanding through multitask learning.
-Entensive Experiments on U-MStance demonstrate that PRISM consistently outperforms strong baselines and exhibits robust generalization.
+Extensive Experiments on U-MStance demonstrate that PRISM consistently outperforms strong baselines and exhibits robust generalization.
+Limitations
+Despite the significant improvements achieved by our model, certain limitations remain.
+First, while U-MStance spans diverse domains including politics, automotive, consumer goods, and cryptocurrency, it lacks coverage of highly specialized or professional sectors, such as specific legal statutes or cutting-edge scientific controversies.
+Second, as illustrated in Figure 5, model performance exhibits a declining trend as conversational depth increases.
+In extremely deep or multi-branch discussion threads, the model still encounters challenges in perfectly capturing subtle logical transitions and long-range contextual drifts.
+Appendix A Detailed Dataset Statistics
+Figure 6 illustrates the distribution of instances across different labels.
+Table 5 illustrates the distribution of instances across different conversational depths, including the number of users and images for each target.
+Within our collected dataset, 36% of the posts contain a single image, while 63% include two or more images, which means visual information is also present within the comments.
+In general, our dataset features more extensive conversational depth compared to existing conversational stance detection datasets, and achieves 100% vision-related coverage.
+Figure 6: The distribution of stance categories in our U-MStance.
+Table 5: Comparison of Depth and Interaction Data Across Different Entities
+Depth Avg. WC Trump Biden Tesla BMW Costco Bitcoin Total
+User 5087 3371 3896 3956 4519 4123 24952
+Image 47 40 55 102 20 21 285
+Post 1 17.96 16 27 14 19 8 13 97 (0.24%)
+Comment 2 27.38 927 1011 723 741 768 1186 5356 (13.39%)
+3 23.71 2306 2096 1982 1816 2288 2647 13135 (32.84%)
+4 27.65 1645 1189 1402 1176 1470 1292 8174 (20.43%)
+5 29.99 1076 694 928 821 1001 846 5366 (13.41%)
+6 31.52 795 427 621 637 541 517 3538 (8.84%)
+7 35.98 529 231 430 385 316 304 2195 (5.49%)
+8 39.59 333 106 264 290 168 173 1334 (3.33%)
+9 45.02 231 65 150 148 108 106 808 (2.02%)
+Appendix B Model Details
+Table 6: Model details including parameters, access type, and their respective sources.
+Model Parameters Access Source
+BERT ¡1B Local Google
+RoBERTa ¡1B Local Meta
+PoliBERTweet ¡1B Local Kawintiranon et al.
+LLaMA-2-7B 7B Local Meta
+GPT-3.5 Unknown API OpenAI
+GPT-4 Unknown API OpenAI
+BERT-ViT ¡1B Local Liang et al.
+TMPT ¡1B Local Liang et al.
+LLaVA-1.5-7B 7B Local Microsoft
+Qwen2.5-VL-7B 7B Local Alibaba
+MiMo-VL-7B 7B Local Xiaomi
+GPT4-1 Unknown API OpenAI
+Appendix C Prompt Templates
+To support the reproducibility and interpretability of our methodology and experimental settings, we provide the comprehensive set of prompt templates used in this work.
+Specifically, Table 7 details the prompts for User Persona Distillation; Table 8 presents the two-stage prompt protocol for Image Intent Extraction; and Table 9 outlines the prompt employed for the baseline models.
+B.1 User Persona Distillation System Prompt: You are an expert in computational psychology and linguistic analysis.
+Your task is to analyze a user’s Big Five personality traits (Openness, Conscientiousness, Extraversion, Agreeableness, and Neuroticism) based on their historical social media activity.
+User Prompt: ### Task Description: Analyze the provided user profiles, including their posts and comments.
+For each of the five traits, provide a qualitative description of how the user exhibits this trait and a quantitative score on a scale of 1-5 (1: Very Low, 5: Very High).
+### Big Five Traits Definition: 1.
+Openness to Experience: Curiosity, originality, and openness to new ideas.
+2.
+Conscientiousness: Organization, responsibility, and goal-oriented behavior.
+3.
+Extraversion: Sociability, assertiveness, and emotional expressiveness.
+4.
+Agreeableness: Trust, altruism, kindness, and affection.
+5.
+Neuroticism: Emotional instability, anxiety, and irritability.
+### Input Data: User ID: {{user id}} Recent Posts: {{post list}} Recent Comments: {{comment list}} Output:
+Table 7: Prompt template for User Persona Distillation. The model outputs a summary of Big Five personality traits for users.
+B.2 Image Intent Extraction Step 1: System Prompt: You are an expert image analyzer.
+User Prompt: ### Task Description: You will be given an image.
+Your task is to produce a detailed and structured description of the image.
+Please follow the required steps: 1.
+**Overall Scene Description** Describe the setting, environment, and general context of the image.
+2.
+**Key Objects and Entities** Identify and describe the notable objects, people, or elements in the image, including their appearance, attributes, and spatial relationships.
+3.
+**Fine-Grained Visual Details** Highlight colors, expressions, textures, gestures, body language, or any small details that may convey meaning.
+4.
+**Style and Emotional Tone (if applicable)** Describe the mood, atmosphere, artistic style, or emotional impression conveyed by the image.
+### Input: ¡image¿ ### Output Requirment: Make sure your description is precise, neutral, and avoids assumptions unless clearly supported by visual evidence.
+Directly output the description text**, starting with **‘Image Description:‘** followed by your concise answer.
+Output: Step 2: System Prompt: You are an expert in multimodal conversation analysis.
+User Prompt: ### Task Description: Determine the **intention or communicative purpose** of the user who posted the image in the conversation.
+Consider the social and semantic context of the conversation, as well as the content of the image.
+You are given: 1.
+A **description of the image**, which has been generated by a visual-language model.
+2.
+A **conversation history** between multiple participants.
+- In this history, the placeholder ”¡image¿” marks where the image was posted by the user.
+If the conversation contains multiple images, carefully identify the **index or order** of the images (e.g., ¡image1¿, ¡image2¿, ¡image3¿…) and determine where the **current target image** appears in the conversation before analyzing it.
+### Requirment: Follow these steps carefully: 1.
+Review the conversation history and understand the flow of dialogue.
+2.
+Interpret the meaning of the image based on the provided description.
+3.
+Analyze how the image contributes to the conversation — for example, whether it expresses emotion, humor, evidence, sarcasm, agreement, mockery, or provides visual information.
+4.
+Infer the **most likely intention** of the user when posting the image.
+5.
+Directly output the intent text**, starting with **‘Intent:‘** followed by your concise answer.
+### Image Description: The image captures a bustling urban scene at night, likely Times Square in New York City, characterized by its iconic billboards and vibrant advertisements.
+The foreground features … ### Conversation History: User youyouxue: We just drove from San Francisco to New York City in a Tesla Model X without using a single drop of gas!
+¡image1¿ User salec1:I did a shorter version of the same thing (Miami to NYC) last month.
+It’s all about … User KickAClay:I always wondered if any Tesla, after inputting your route … User timdorr:Yes, it does exactly this.
+Set your destination and it will find Superchargers along your route, tell you how long to … User chupacabrando:How much longer does it take … Output:
+Table 8: Prompt template for Image Intent Extraction. The model outputs a detailed and structured description of the image.
+B.3 Zero-shot Stance Detection System Prompt: You are an expert linguist specializing in conversational analysis and social media stance detection.
+User Prompt: ### Task Description: Analyze the provided conversation flow and determine the stance of the current speaker toward the specific Target listed below.
+The attached image(s) correspond to the sequential labels (e.g., ¡image1¿, ¡image2¿) mentioned in the text.
+### Target: Tesla ### Stance Classes: [’favor’,’against’,’none’] ### Conversation History: User youyouxue: We just drove from San Francisco to New York City in a Tesla Model X without using a single drop of gas!
+¡image1¿ User salec1:I did a shorter version of the same thing (Miami to NYC) last month.
+It’s all about … User KickAClay:I always wondered if any Tesla, after inputting your route … User timdorr:Yes, it does exactly this.
+Set your destination and it will find Superchargers along your route, tell you how long to … User chupacabrando:How much longer does it take … ### Current Speaker’s Reddit Profile: Active Communities: — Historical Posts — — Historical Comments — ### Output Format (JSON): s̈tance:̈ ”favor” — ”against” — ”none” Output:
+Table 9: Prompt template for Multimodal Conversational Stance Detection. The model outputs a stance class for the current instance.